The height of a triangle is six feet less than twice the base. If the area is 70 squae feet. Find the base and the height
Answer by checkley77(12844)

H=2B-6
A=BH/2
70=B(2B-6)/2
70=(2B^2-6B)/2
2B^2-6B=2*70
2B^2-6B=140
2B^2-6B-140=0
2(B^2-3B-70)=0
2(B-10)(B+7)=0
B-10=0
B=10 ANS. FOR THE BASE.
H=2*10-6=20-6=14 ANS. FOR THE HEIGHT.
PROOF:
70=10*14/2
70=140/2
70=70